OpenStudy (michellemn):

Generate the first 5 terms of this sequence: f(1) = 2 and f(2) = 3, f(n) = f(n - 1) + f(n - 2), for n > 2. 2, 3, 5, 7, 9 2, 3, 4, 7, 11 2, 3, 5, 9, 11 2, 3, 5, 8, 13

OpenStudy (anonymous):

For: n>2 : f=the previous + the one before the previous so for n = 3 f = 2+3 = 5

OpenStudy (anonymous):

for n = 4 f = 3 + 5 = 8

OpenStudy (anonymous):

What are you suppose to be Michelle Obama????? So you the one thats been screwing up our lunch menus!!!!

OpenStudy (anonymous):

Ain;t got nothin to say now do you

OpenStudy (michellemn):

Can you help me with another one? @114Bytes Use the graph below to fill in the blank with the correct number: f(0) = _______ X, Y graph. X range is negative 6 to 6, and Y range is negative 4 to 4. Plotted points negative 3, 0; negative 2, 2; 0, 1; and 1, negative 2. Numerical Answers Expected! Answer for Blank 1:

OpenStudy (anonymous):

o.O What was that all about @jimmiedeel

OpenStudy (michellemn):

I have no idea. I read his post and I got really confused.

OpenStudy (anonymous):

Give me a minute to figure this one out, my approach is that i see you want the value f(0) assuming y is a function of x and we're plugging x = 0 into the equation So first step to find the equation, if it's a line it's easy. y-yo=m(x-xo)

OpenStudy (michellemn):

ok

OpenStudy (anonymous):

f(0) = 1 Because y why is a function of x so we can say f(x) = y = some function with x in it

OpenStudy (anonymous):

So when x = 0 y = ?

OpenStudy (anonymous):

(0,1)

OpenStudy (michellemn):

So the answer is: (0,1)

OpenStudy (anonymous):

If the plotted points are (x,y) then go ahead and tell me what y is. when x=0

OpenStudy (michellemn):

1

OpenStudy (anonymous):

:)

OpenStudy (michellemn):

so f(0) = 1

OpenStudy (michellemn):

thanks

OpenStudy (anonymous):

No problem! Let me know if you have any confusion :)